OpenStudy (anonymous):

Which of the following is not a Pythagorean triple? A. 28, 45, 53 B. 16, 63, 65 C. 13, 84, 85 D. 11, 61, 62

jimthompson5910 (jim_thompson5910):

Something like 3,4,5 is a Pythagorean triple because \[\Large 3^2 + 4^2 = 5^2\] is a true equation

OpenStudy (richyw):

Check to see if the points satisfy\[a^2+b^2=c^2\]with \(c>b,a\)

jimthompson5910 (jim_thompson5910):

But something like 1,2,3 is NOT a Pythagorean triple because \[\Large 1^2 + 2^2 = 3^2\] is false

OpenStudy (anonymous):

I got that they were all true..

jimthompson5910 (jim_thompson5910):

one of them is false though, so double check your work

OpenStudy (anonymous):

is D wrong? bc its comes out to be 62.3698...

OpenStudy (richyw):

yes D the correct answer. I always hated these questions! protip: if you can't immediately eliminate easy ones, start from D, teachers love to make them the last one! this could be two true-false questions!

OpenStudy (anonymous):

okay. thank you so much